A merchant has a beaker containing 24 ounces of a precious fluid. He also has empty 5-ounce, 11-
ounce, and 13-ounce beakers. How can he divide the fluid into three equal portions?
1
Expert's answer
2011-10-07T11:33:26-0400
Here is an algoritm
1) full 13-ounce beaker
[table]

5-ounce

11-ounce

13-ounce



0

0

13

[/table]
2)from 13-ounce fill the 11-ounce beaker(at 13-oun. Wewill have 2 ounce of water)
[table]

5-ounce

11-ounce

13-ounce





11

2

[/table]
3)empty 11-oun. And fill it from 13 oun. Wich has 2 ounce of w.
[table]

5-ounce

11-ounce

13-ounce





2



[/table]
4)fill 13-oun. And pour water from it to 5-ouncebeaker
[table]

5-ounce

11-ounce

13-ounce



5

2

13-5=8(!!!)

[/table]
5)
Empty 5-ounce and fill it from 11-ounce
[table]

5-ounce

11-ounce

13-ounce



2

0

8

[/table]6)
Fill 11-ounce and add water to 5-ouncebeaker
[table]

5-ounce

11-ounce

13-ounce



5

11-3=8

8

[/table]7)
Return 5ounce to 24-ounce beaker and itwill have also 5+3=8 ounce of water